# Selective Passive Tuning of Cavity Resonance by Mode Index Engineering of the Partial Length of a Cavity

### v1

@@
Cavities in large-scale photonic integrated circuits often suffer from a wider distribution of resonance frequencies due to fabrication errors. It is crucial to adjust the resonances of cavities using post-processing methods to minimize the frequency distribution. We have developed a concept of passive tuning by manipulating the mode index of a portion of a microring cavity. Through analytical studies and numerical experiments, we have found that depositing a thin film of dielectric material on top of the cavity or etching the material enables us to fine-tune the resonances and minimize the frequency distribution. This versatile method allows for the selective tuning of each cavity’s resonance in a large set of cavities in a single fabrication step, providing robust passive tuning in large-scale photonic integrated circuits. Furthermore, this method can be applied and explored in various cavities and different material configurations.
@@ 1 Introduction
On-chip microring resonators are widely used in various fields due to their small size, high-quality resonances, and advantages in precise manipulation of light [1]. In telecommunications, they enable wavelength division multiplexing (WDM) and modulation of light for increased data transmission over optical fibers [2]. In data centers, they are used in optical interconnects for high-speed and energy-efficient communication [3, 4]. They are also crucial in biosensing for label-free detection of biomolecules and in environmental monitoring to detect trace chemicals [5]. In quantum computing, they are used to generate entangled photon pairs and single-photon sources [6, 7, 8], while in LIDAR systems, they improve autonomous vehicle navigation through frequency comb generation [9, 10]. Microring resonators are integral to on-chip spectrometers [11, 12], optical signal processing [13], and biomedical imaging [14], particularly in Optical Coherence Tomography (OCT) [15]. They are also used in neuromorphic computing for AI [16], atomic clock stabilization [17], and the generation of supercontinuum light sources [18, 19]. In astrophotonics, they assist in exoplanet detection [20], and in optical computing and photonic neural networks, they form reliable fundamental elements of chip [21, 22]. Moreover, microring resonators are used in terahertz generation [23, 24] and secure photonic communication [25]. They also facilitate RF signal processing in microwave photonics, soil, and crop monitoring in agriculture [26, 27], and precision measurements in optical frequency metrology [28]. These applications demonstrate the broad impact of microring resonators and PICs across industries, from healthcare and environmental monitoring to advanced computing and telecommunications.
@@ 2 Resonance shift due to fabrication errors in $L$ and $n_{eff}$
$\frac{\delta f}{f}=\frac{f^{\prime}-f}{f}=\frac{-\delta L}{L+\delta L}$ (4)
@@
$\frac{\delta f}{f}=\frac{f^{\prime}-f}{f}=\frac{-\delta n_{eff}}{n_{eff}+\delta n_{eff}}$ (6)
@@ 3 Partial length mode index $n_{eff}$ manipulation
where $\delta n=n^{\prime}_{eff}-n_{eff}$. When $\delta n$ is negative, the new frequency is higher than the original frequency, i.e., resonance is blueshift, and when $\delta n$ is positive, the new frequency is lower than the original frequency, i.e., resonance is redshift. Using the above equation, we can evaluate the relative shift in resonance frequency when we change the effective mode index in the partial length $L^{\prime}$ of a cavity. Fig. 2 shows the absolute relative change in frequency due to deposition of dielectric material on length $L^{\prime}$. The smaller the change in index $\delta n$ and length ratio $L^{\prime}/L$, the smaller the shift in resonance is achievable and vice versa.
@@ 4 Passive Tuning of cavity’s resonance Selectively through Mode Index Engineering
We exploit the concept discussed above that the shift in resonance frequency of a microring cavity is achievable by introducing a change in the effective index of mode by thin film deposition of a dielectric material on top of the cavity (Fig. 1). We have the freedom to choose the length $L^{\prime}$, i.e., the partial length of the microring cavity, to selectively manipulate the effective index of mode; we define this method of engineering the total optical path length as mode index engineering. Assuming that the cavities are open to air without any cladding or have minimal cladding on top such that if there is an introduction of dielectric material on top of the cavity, then it could achieve a change in the effective index of the mode significantly within a desired range of resonance shift, the tuning range and resolution can be understood from Fig. 2. Fig. 3 shows a schematic diagram of a set of cavities tuned selectively by the proposed idea of mode index manipulation of partial length. Due to the inherent scalability of this method, the tuning operation can be executed in a single fabrication process.
Figure 3: Tuning the resonances of microring cavities selectively by individually selecting partial microring cavity length of each cavity. A set of one type cavities in an array of size m x k on a chip. The length ($L^{\prime}$) is evaluated by the difference between the original frequency and the desired frequency, and detailed numerical analysis of $n_{eff}$, $\delta n$, and $\delta f(L^{\prime})/f$ relation. The process can also be implemented on multiple chips with multiple cavities, thus enabling us to reduce frequency distribution from chip to chip in a large-scale manufacturing process.
@@ 4.2 Passive tuning of a silicon nitride resonator resonance
We discuss an example of silicon nitride ($Si_{3}N_{4}$) microring resonator and apply numerical analysis for red-shifting the resonance frequency. We use finite-difference time-domain (FDTD) analysis in our numerical experiments. In complementary metal-oxide-semiconductor (CMOS) technology, dielectric materials like silicon dioxide ($SiO_{2}$), aluminum oxide ($Al_{2}O_{3}$), and silicon oxynitride (SiON) are commonly used in photonic integrated circuits. We choose two configurations (A and B) for numerical demonstration purposes, as shown in Fig. 4. The first case (A) involves the original configuration of the microring resonator core waveguide being open to the air, while the second case (B) includes a cladding deposited on the microring resonator core waveguide layer structure. We estimate the change in the effective index of fundamental $TE_{0}$ mode due to the deposition of a thin film dielectric material and investigate the tunable range and resolution achievable by our method. Note that this method is applicable to any higher confined mode in a cavity. We consider $Si_{3}N_{4}$ as core waveguide material and $SiO_{2}$ and $Al_{2}O_{3}$ as cladding materials. We assume the refractive index of $Si_{3}N_{4}$ is 2, $Al_{2}O_{3}$ is 1.76, and $SiO_{2}$ is 1.457 at 772.5 nm. Also, note that the material and geometry dispersion are included in the estimation of effective mode index $n_{eff}$. Therefore, the method applies to dispersive materials and accounts for material and geometric dispersions.
@@
Figure 4: The schematics diagram of the layer structure cross-sections of two configurations. First configuration (A): a) cavity core waveguide material $Si_{3}N_{4}$ (yellow color) on top of $SiO_{2}$ cladding (grey color). b) A dielectric material layer (dark grey color) is deposited on the partial length of the cavity to tune the cavity’s resonance selectively (see Fig. 1 and 3). Second configuration (B): c) cavity core waveguide material $Si_{3}N_{4}$ (yellow color) on top of $SiO_{2}$ cladding (grey color) with a cladding (green color) on top of $Si_{3}N_{4}$. d) A dielectric material layer (dark grey color) is deposited on a partial length of the cavity to tune the cavity’s resonance selectively. In part c), the top cladding (green color) thickness should be small enough to enable the desired post-processing selective tuning of each cavity’s resonance.
@@
Figure 5: Mode profile (|E(x,y)|) of $TE_{0}$ mode confined in 100 nm thick 4000 nm wide $Si_{3}N_{4}$ core waveguide for three cases: a) on $SiO_{2}$ cladding (see Fig. 4(a)), b) on $SiO_{2}$ cladding with $SiO_{2}$ thin film material of thickness 10 nm deposited on top (see Fig. 4(b)), and c) on $SiO_{2}$ cladding with $Al_{2}O_{3}$ thin film material of thickness 10 nm deposited on top (see Fig. 4(b)). White and black dotted lines show $Si_{3}N_{4}$ core waveguide and lower $SiO_{2}$ cladding interface with $Si_{3}N_{4}$, respectively.
@@
Figure 6: The plot of change in effective mode index ($\delta n$) of $TE_{0}$ mode confined in 100 nm thick 4 $\mu$m wide $Si_{3}N_{4}$ core waveguide with varied thickness of deposited dielectric materials $SiO_{2}$ and $Al_{2}O_{3}$ (configuration A, see Fig. 4(b)): a) 0-11 nm thickness range on x-axis and b) 0-380 nm thickness range on x-axis.
@@
In the numerical experiment, we assume a 100 nm thick 4000 nm wide $Si_{3}N_{4}$ core waveguide on top of $SiO_{2}$ cladding for the microring resonator. Fig. 5(a) shows the mode profile of fundamental TE (TE0) mode confined in the core waveguide, Fig. 5(b) shows the mode profile of TE0 mode confined in $Si_{3}N_{4}$ core waveguide with 10 nm $SiO_{2}$ deposited on top and Fig. 5(c) shows the mode profile of TE0 mode confined in $Si_{3}N_{4}$ core waveguide with 10 nm $Al_{2}O_{3}$ deposited on top. Since the refractive index of $Al_{2}O_{3}$ is larger than $SiO_{2}$, the change in the effective index of mode is larger for a given thickness, as shown in Fig. 6. Even for a small thickness $\sim$ 1 nm of deposited material (see Fig. 6 (a)), the change in the effective index, $\delta n$ is $\sim$ 0.001. In configuration B (see Fig. 4(c)), a certain thickness of cladding is on top of the core waveguide already, and the original resonance frequency is defined for this configuration B (Fig. 4(c)). The tuning step is performed by depositing a dielectric material (see Fig. 4(d)). As illustrated in Fig. 6 and 7(a), the curve saturates for large thickness, and $\delta n$ can be an order of $\sim 0.0001$ for the configuration B, which is suitable for ultra-fine tuning purposes. Fig. 7(b) shows the change in the effective index of mode due to $Al_{2}O_{3}$ material deposition on a $SiO_{2}$ cladding of thickness 300 nm or 400 nm on top of the core waveguide. The change in $\delta n$ due to $Al_{2}O_{3}$ is small for 400 nm $SiO_{2}$ cladding compared to the 300 nm case, which is obvious due to the confinement of mode in the core waveguide and decrease in mode overlap. We can learn from this analysis and include the role of cladding thickness to optimize $\delta n$ for our specific application and needs.
@@
The second configuration (B) is described in Fig. 4(c), so the original frequency of the cavity’s resonance is assumed for this configuration, and the desired frequency is set to this value for their application in PICs. The mode is confined in the core waveguide and the mode overlap with the top cladding is small. Adding another thin dielectric material on top of this cladding would produce a very small shift in the effective index of mode compared to the cavity without cladding on top. Fig. 6(b) and 7(a) shows the curve for change in the effective mode index with the cladding thickness deposited on top of 100 thick 4000 nm wide $Si_{3}N_{4}$ waveguide. The curve reaches saturation after a certain thickness. Before the saturation, the resonance is less sensitive. This region near saturation allows the ultra-fine tune of the cavity’s resonance by choosing $\delta n\sim 0.0001$ or even smaller value appropriately.
Figure 7: Plot of change in effective mode index ($\delta n$) of $TE_{0}$ mode confined in 100 nm thick 4 $\mu$m wide $Si_{3}N_{4}$ core waveguide: a) due to deposition of $SiO_{2}$ dielectric material on top of $Si_{3}N_{4}$, $Si_{3}N_{4}$ sits on top of $SiO_{2}$ cladding (configuration A, see Fig. 4(b)). The inset shows a small range of data (red marked area); the change in $\delta n$ is achievable as small as $\sim 0.0001$ for even large $\sim$ 50 nm thick material deposition if configuration B (see Fig. 4 (c) and (d)) is implemented, b) due to deposition of $Al_{2}O_{3}$ on top of 300 nm (blue dotted scatter points) and 400 nm (red dotted scatter points) thick $SiO_{2}$ cladding on $Si_{3}N_{4}$, $Si_{3}N_{4}$ sits on top of $SiO_{2}$ cladding (configuration B, see Fig. 4 (c) and (d)). In this case, $\delta n\sim$ 0.0001 is attainable for about 5-10 nm thick $Al_{2}O_{3}$. The thickness of $SiO_{2}$ cladding can be optimized (see Fig. 4(c)) based on the desired $\delta n$ in the tuning operation.

### v2

@@
Cavities in large-scale photonic integrated circuits often suffer from a wider distribution of resonance frequencies due to fabrication errors. It is crucial to adjust the resonances of cavities using post-processing methods to minimize the frequency distribution. We have developed a concept of passive tuning by manipulating the mode index of a portion of a microring cavity. Through analytical studies and numerical experiments, we have found that depositing a thin film of dielectric material on top of the cavity or etching the material enables us to fine-tune the resonances and minimize the frequency distribution. This versatile method allows for the selective tuning of each cavity’s resonance in a large set of cavities in a single fabrication step, providing robust passive tuning in large-scale photonic integrated circuits. We show that proposed method achieves tuning resolution below 1/Q and range upto $10^{3}/Q$ for visible to near-infrared wavelengths. Furthermore, this method can be applied and explored in various optical cavities and material configurations.
@@ 1 Introduction
On-chip microring resonators are widely used in various fields due to their small size, high-quality resonances, and advantages in precise manipulation of light [1]. In telecommunications, they enable wavelength division multiplexing (WDM) and modulation of light for increased data transmission over optical fibers [2]. In data centers, they are used in optical interconnects for high-speed and energy-efficient communication [3, 4]. They are also crucial in biosensing for label-free detection of biomolecules and in environmental monitoring to detect trace chemicals [5]. In quantum computing, they are used to generate entangled photon pairs and single-photon sources [6, 7, 8], while in LIDAR systems, they improve autonomous vehicle navigation through frequency comb generation [9, 10]. Microring resonators are integral to on-chip spectrometers [11, 12], optical signal processing [13], and biomedical imaging [14], particularly in Optical Coherence Tomography (OCT) [15]. They are also used in neuromorphic computing for AI [16], atomic clock stabilization [17], and the generation of supercontinuum light sources [18, 19]. In astrophotonics, they assist in exoplanet detection [20], and in optical computing and photonic neural networks, they form reliable fundamental elements of chip [21, 22]. Moreover, microring resonators are used in terahertz generation [23, 24] and secure photonic communication [25]. They also facilitate RF signal processing in microwave photonics, soil, and crop monitoring in agriculture [26, 27], and precision measurements in optical frequency metrology [28]. Coupled-resonator optical waveguide (CROW) devices have advantages in development of complex photonic devices for non-linear photonics, optical band-pass filters, chip-integrated lasers, and sensors. These applications demonstrate the broad impact of microring resonators and PICs across industries, from healthcare and environmental monitoring to advanced computing and telecommunications.
@@ 2 Resonance shift due to fabrication errors in $L$ and $n_{eff}$
$\frac{\delta f}{f}=\frac{f^{\prime}-f}{f}=\frac{-\delta L}{L+\delta L}\sim\frac{-\delta L}{L}$ (4)
@@
$\frac{\delta f}{f}=\frac{f^{\prime}-f}{f}=\frac{-\delta n_{eff}}{n_{eff}+\delta n_{eff}}\sim\frac{-\delta n_{eff}}{n_{eff}}$ (6)
@@ 3 Partial length mode index $n_{eff}$ manipulation
where $\delta n=n^{\prime}_{eff}-n_{eff}$. When $\delta n$ is negative, the new frequency is higher than the original frequency, i.e., resonance is blueshift, and when $\delta n$ is positive, the new frequency is lower than the original frequency, i.e., resonance is redshift. For small values of $\delta n$, the eq. 8 can be simplified to, $\frac{\delta f}{f}\sim-\frac{L^{\prime}\delta n}{Ln_{eff}}$. Using eq. 8, we can evaluate the relative shift in resonance frequency when we change the effective mode index in the partial length $L^{\prime}$ of a cavity. Fig. 2 shows the absolute relative change in frequency due to deposition of dielectric material on length $L^{\prime}$. The smaller the change in index $\delta n$ and length ratio $L^{\prime}/L$, the smaller the shift in resonance is achievable and vice versa.
@@ 4 Passive Tuning of cavity’s resonance Selectively through Mode Index Engineering
We exploit the concept discussed above that the shift in resonance frequency of a microring cavity is achievable by introducing a change in the effective index of mode by thin film deposition of a dielectric material on top of the cavity (Fig. 1). We have the freedom to choose the length $L^{\prime}$, i.e., the partial length of the microring cavity, to manipulate the effective index of mode selectively; we define this method of engineering the total optical path length as mode index engineering. Assuming that the cavities are open to air without any cladding or have minimal cladding on top such that if there is an introduction of dielectric material on top of the cavity, then it could achieve a change in the effective index of the mode significantly within a desired range of resonance shift, the tuning range and resolution can be understood from Fig. 2. Fig. 3 shows a schematic diagram of a set of cavities tuned selectively by the proposed idea of mode index manipulation of partial length. Due to the inherent scalability of this method, the tuning operation can be executed in a single fabrication process.
Figure 3: Tuning the resonances of microring cavities selectively by individually selecting partial microring cavity length of each cavity. A set of one type cavities in an array of size m $\times$ k on a chip. The length ($L^{\prime}$) is evaluated by the difference between the original frequency and the desired frequency, and detailed numerical analysis of $n_{eff}$, $\delta n$, and $\delta f(L^{\prime})/f$ relation. The process can also be implemented on multiple chips with multiple cavities, thus enabling us to reduce frequency distribution from chip to chip in a large-scale manufacturing process.
@@ 4.2 Passive tuning of a silicon nitride resonator resonance
We discuss an example of silicon nitride ($Si_{3}N_{4}$) microring resonator and apply numerical analysis for red-shifting the resonance frequency. We use finite-difference time-domain (FDTD) analysis in our numerical experiments. In complementary metal-oxide-semiconductor (CMOS) technology, dielectric materials like silicon dioxide ($\text{SiO}_{2}$), aluminum oxide ($\text{Al}_{2}\text{O}_{3}$), and silicon oxynitride (SiON) are commonly used in photonic integrated circuits. We choose two configurations (A and B) for numerical demonstration purposes, as shown in Fig. 4. The first case (A) involves the original configuration of the microring resonator core waveguide being open to the air, while the second case (B) includes a cladding deposited on the microring resonator core waveguide layer structure. We estimate the change in the effective index of fundamental $TE_{0}$ mode due to the deposition of a thin film dielectric material and investigate the tunable range and resolution achievable by our method. Note that this method is applicable to any higher confined mode in a cavity. We consider $\text{Si}_{3}\text{N}_{4}$ as core waveguide material and $\text{SiO}_{2}$ and $\text{Al}_{2}\text{O}_{3}$ as cladding materials. We assume the refractive index of $\text{Si}_{3}\text{N}_{4}$ is 2, $\text{Al}_{2}\text{O}_{3}$ is 1.76, and $\text{SiO}_{2}$ is 1.457 at 772.5 nm. Also, note that the material and geometry dispersion are included in the estimation of effective mode index $n_{eff}$. Therefore, the method applies to dispersive materials and accounts for material and geometric dispersions.
@@
Figure 4: The schematic diagram of the layer structure cross-sections of two configurations. First configuration (A): a) cavity core waveguide material $\text{Si}_{3}\text{N}_{4}$ (yellow color) on top of $\text{SiO}_{2}$ cladding (grey color). b) A dielectric material layer (dark grey color) is deposited on the partial length of the cavity to tune the cavity’s resonance selectively (see Fig. 1 and 3). Second configuration (B): c) cavity core waveguide material $\text{Si}_{3}\text{N}_{4}$ (yellow color) on top of $\text{SiO}_{2}$ cladding (grey color) with a cladding (green color) on top of $\text{Si}_{3}\text{N}_{4}$. d) A dielectric material layer (dark grey color) is deposited on a partial length of the cavity to tune the cavity’s resonance selectively. In part c), the top cladding (green color) thickness should be small enough to enable the desired post-processing selective tuning of each cavity’s resonance.
@@
Figure 5: Mode profile (|E(x,y)|) of $TE_{0}$ mode confined in 100 nm thick 4000 nm wide $\text{Si}_{3}\text{N}_{4}$ core waveguide for three cases: a) on $\text{SiO}_{2}$ cladding (see Fig. 4(a)), b) on $\text{SiO}_{2}$ cladding with $\text{SiO}_{2}$ thin film material of thickness 10 nm deposited on top (see Fig. 4(b)), and c) on $\text{SiO}_{2}$ cladding with $\text{Al}_{2}\text{O}_{3}$ thin film material of thickness 10 nm deposited on top (see Fig. 4(b)). White and black dotted lines show $\text{Si}_{3}\text{N}_{4}$ core waveguide and lower $\text{SiO}_{2}$ cladding interface with $Si_{3}N_{4}$, respectively.
@@
Figure 6: The plot of change in effective mode index ($\delta n$) of $TE_{0}$ mode confined in 100 nm thick 4 $\mu$m wide $\text{Si}_{3}\text{N}_{4}$ core waveguide with varied thickness of deposited dielectric materials $\text{SiO}_{2}$ and $\text{Al}_{2}\text{O}_{3}$ (configuration A, see Fig. 4(b)): a) 0-11 nm thickness range on x-axis and b) 0-380 nm thickness range on x-axis.
@@
In the numerical experiment, we assume a 100 nm thick 4000 nm wide $\text{Si}_{3}\text{N}_{4}$ core waveguide on top of $\text{SiO}_{2}$ cladding (lower cladding) for the microring resonator. Fig. 5(a) shows the mode profile of fundamental TE (TE0) mode confined in the core waveguide, Fig. 5(b) shows the mode profile of TE0 mode confined in $\text{Si}_{3}\text{N}_{4}$ core waveguide with 10 nm $\text{SiO}_{2}$ deposited on top and Fig. 5(c) shows the mode profile of TE0 mode confined in $\text{Si}_{3}\text{N}_{4}$ core waveguide with 10 nm $\text{Al}_{2}\text{O}_{3}$ deposited on top. Since the refractive index of $\text{Al}_{2}\text{O}_{3}$ is larger than $\text{SiO}_{2}$, the change in the effective index of mode is larger for a given thickness, as shown in Fig. 6. Even for a small thickness $\sim$ 1 nm of deposited material (see Fig. 6 (a)), the change in the effective index, $\delta n$ is $\sim$ 0.001. In configuration B (see Fig. 4(c)), a certain thickness of cladding is on top of the core waveguide already, and the original resonance frequency is defined for this configuration B (Fig. 4(c)). The tuning step is performed by depositing a dielectric material (see Fig. 4(d)). As illustrated in Fig. 6 and 7(a), the curve saturates for large thickness, and $\delta n$ can be an order of $\sim 0.0001$ for the configuration B, which is suitable for ultra-fine tuning purposes. Fig. 7(b) shows the change in the effective index of mode due to $\text{Al}_{2}\text{O}_{3}$ material deposition on a $\text{SiO}_{2}$ cladding of thickness 300 nm or 400 nm on top of the core waveguide. The change in $\delta n$ due to $\text{Al}_{2}\text{O}_{3}$ is small for 400 nm $\text{SiO}_{2}$ cladding compared to the 300 nm case, which is obvious due to the confinement of mode in the core waveguide and decrease in mode overlap. We can learn from this analysis and include the role of mid cladding thickness to optimize $\delta n$ for our specific application and needs.
@@
The second configuration (B) is described in Fig. 4(c), so the original frequency of the cavity’s resonance is assumed for this configuration, and the desired frequency is set to this value for their application in PICs. The mode is confined in the core waveguide and the mode overlap with the top cladding is small. Adding another thin dielectric material on top of this cladding would produce a very small shift in the effective index of mode compared to the cavity without cladding on top. Fig. 6(b) and 7(a) shows the curve for change in the effective mode index with the cladding thickness deposited on top of 100 thick 4000 nm wide $\text{Si}_{3}\text{N}_{4}$ waveguide. The curve reaches saturation after a certain thickness. Before the saturation, the resonance is less sensitive. This region near saturation allows the ultra-fine tune of the cavity’s resonance by choosing $\delta n\sim 0.0001$ or even smaller value appropriately.
Figure 7: Plot of change in effective mode index ($\delta n$) of $TE_{0}$ mode confined in 100 nm thick 4 $\mu$m wide $\text{Si}_{3}\text{N}_{4}$ core waveguide: a) due to deposition of $\text{SiO}_{2}$ dielectric material on top of $Si_{3}N_{4}$, $\text{Si}_{3}\text{N}_{4}$ sits on top of $\text{SiO}_{2}$ cladding (configuration A, see Fig. 4(b)). The inset shows a small range of data (red marked area); the change in $\delta n$ is achievable as small as $\sim 0.0001$ for even large $\sim$ 50 nm thick material deposition if configuration B (see Fig. 4 (c) and (d)) is implemented, b) due to deposition of $\text{Al}_{2}\text{O}_{3}$ on top of 300 nm (blue dotted scatter points) and 400 nm (red dotted scatter points) thick $\text{SiO}_{2}$ cladding on $Si_{3}N_{4}$, $\text{Si}_{3}\text{N}_{4}$ sits on top of $\text{SiO}_{2}$ cladding (configuration B, see Fig. 4 (c) and (d)). In this case, $\delta n\sim$ 0.0001 is attainable for about 5-10 nm thick $\text{Al}_{2}\text{O}_{3}$. The thickness of $\text{SiO}_{2}$ cladding can be optimized (see Fig. 4(c)) based on the desired $\delta n$ in the tuning operation.